# Looking Backward: Retrospective Backward Synthesis for Goal-Conditioned GFlowNets

### v1

@@ 4.3 Proposed Method
In this section, we propose a novel approach, Retrospective Backward Synthesis (RBS), which is a simple yet effective method to efficiently tackle this challenging problem in GC-GFlowNets.
@@
Consider a trajectory $\tau=\{s_{0}\to\cdots\to s_{i}\to\cdots\to s_{n}\}$ collected by the forward policy $P_{F}$ of GC-GFlowNets which fails to reach the goal ($s_{n}\neq y$) with a zero reward. As illustrated in Fig. 2, RBS utilizes the potential of the backward policy $P_{B}$ to synthesize a reverse trajectory $\tau^{\prime}=\{y\to\cdots\to s_{i}^{\prime}\to\cdots\to s_{0}\}$ from the commanded goal. Therefore, $\tau^{\prime}$ provides a successful training experience as it achieves the goal state, thus enriching training data with positive feedback for GC-GFlowNets to mitigate the sparse reward problem. Unlike previous reward relabeling techniques, such as HER (Andrychowicz et al., 2017), which simply replaces the original goal $y$ with the achieved state $s_{n}$, RBS provides more diverse and informative new training data by imaging a new trajectory $\tau^{\prime}\neq\tau$, thus leading to more sample-efficient learning. In practice, we store both collected trajectories $\{\tau_{i}\}_{i=1}^{m}$ and experiences from RBS $\{\tau^{{}^{\prime}}_{i}\}_{i=1}^{m}$ in the replay buffer, and jointly replay the two types of trajectories to optimize GC-GFlowNets.
Next, we introduce the techniques we develop to guarantee the efficiency and effectiveness of training GC-GFlowNets with RBS, where the overall algorithm is summarized in Alg. 1.
@@
3:     Sample a goal $y$ randomly
@@
Age-Based Sampling. To optimize data utilization, we leverage a replay buffer (Lin, 1992, Schaul et al., 2016) to store both the trajectories collected by GC-GFlowNets and the retrospective backward synthesized experiences. To guarantee that all experiences are fully considered during training, we introduce an age-based sampling technique. Specifically, age-based sampling assigns the highest priority $p_{\rm max}>0$ to newly added experiences and updates their priority to zero after being learned. Consequently, newly added experiences can be replayed first, while learned experiences are randomly sampled. This prioritization scheme ensures that experiences are leveraged more thoroughly, which balances the exploration of fresh experiences and the exploitation of acquired knowledge.
Backward Policy Regularization. The backward policy can be chosen freely as studied in (Bengio et al., 2023). In the extreme case where $P_{B}$ is set to be a uniform policy, the optimization of GC-GFlowNets becomes challenging as it is not learnable and the data is excessively diverse. On the other hand, specifying $P_{B}$ to be a deterministic policy can limit the data diversity. We therefore learn the backward policy $P_{B}(\cdot|s^{\prime},\theta)$ within GC-GFlowNets based on the flow consistency criterion in §4.1, instead of specifying it to be a fixed policy, which offers smooth sampling aligned with the current model’s capacity. Yet, when $P_{B}$ degenerates into a deterministic policy, it fails to provide diverse backward trajectories, which can limit the potential of GC-GFlowNets to generalize well. To strike a balance between these two extremes and further enhance the diversity of the imagined trajectories, we introduce a backward policy regularization for $P_{B}$. This regularization term penalizes the Kullback-Leibler (KL) divergence between $P_{B}(\cdot|s^{\prime},\theta)$ and a uniform distribution $\mathcal{U}$, thus encouraging $P_{B}$ to resemble a uniform distribution, while allowing for learning and adaptation. Our training objective can be written as in Eq. (6), where $\gamma$ is the regularization coefficient.
@@
Intensified Reward Feedback. For long-horizon and high-dimensional tasks, a critical factor that affects learning effectiveness is the efficient propagation of the reward signal, which may require a number of steps and affect the learning of intermediate steps. We propose an efficient technique that intensifies the learning signal, defined as $\mathcal{L}_{\rm GC-GFN}=\left(\log\left[{F_{\theta}(s|y)P_{F}(s^{\prime}|s,y,\theta)}\right]-\log\left[{CR(x,y)P_{B}(s|s^{\prime},y,\theta)}\right]\right)^{2}$ for terminal states $s^{\prime}$, where $C$ is a intensification coefficient to scale the effect of $R(x,y)$. The mechanism behind this technique is that a larger value of $C$ indeed amplifies the gradient of $P_{B}$ by $\log(C\times R(x,y))$ for terminal states (the detailed derivation can be found in Appendix A). By the intensified reward feedback with a large value of $C$, we effectively strengthen the learning signal propagated backward through the trajectory, enabling more efficient learning and faster convergence in complex environments.
@@
Empirical Validation Fig. 1 compares RBS and GC-GFN with HER (Andrychowicz et al., 2017) in the set generation task with increasing set sizes. The result demonstrates that our RBS approach can scale up to problems with large set sizes, while GC-GFlowNet with HER fails to maintain good performance as the problem complexity increases.
@@ 5 Experiments
In this section, we conduct extensive experiments to investigate our Retrospective Backward Synthesis (RBS) method to answer the following key questions:
  • How does RBS-GFN compare with baselines in terms of sample efficiency and success rates?
  • Can RBS-GFN scale to more complex environments?
  • Can RBS-GFN effectively generalize to unseen goals and varying environments?
  • Can RBS-GFN be built upon different GFlowNets training objectives?
  • What’s the influence of employing different backward policies for backward trajectory synthesis?
@@ 5.1 GridWorld
In the GridWorld environment (Bengio et al., 2021), the model learns to achieve any given goals starting in a $H\times H$ grid. Specifically, we investigate mazes with increasing horizons $H$ ($32$, $64$, and $128$), respectively, resulting in different levels of difficulty categorized as small, medium, and large.
We compare our proposed RBS-GFN approach with the following state-of-the-art baselines. (i) GFN w/ HER (Andrychowicz et al., 2017) is a GC-GFlowNets that relabels the negative reward in a failed trajectory with a positive reward. (ii) OC-GAFN (Pan et al., 2023a) is a recent method that utilizes contrastive learning to complement successful experiences, and employs a trained Generative Augmented Flow Network (GAFN; Pan et al. (2023c)) as an exploratory component to generate diverse outcomes $y$, which are subsequently provided to sample goal-conditioned trajectories. (iii) DQN w/ HER leverages both deep Q-learning algorithm (Mnih et al., 2013, Jang et al., 2019) and HER technique (Andrychowicz et al., 2017) to learn a near-optimal policy. This baseline is used to ablate the effects of GFlowNet-based training compared with RL style methods. To ensure fairness, each baseline has the same model architecture and training steps as RBS-GFN, and we follow the experimental setup for hyperparameters as in (Pan et al., 2023a). We run each algorithm with three different seeds and report their performance in mean and standard deviation. A more detailed description of the experimental setup can be found in Appendix C.
@@ 5.1.1 Performance Comparison
The success rates for different methods for increasing sizes of the GridWorld environment (including small, medium, and large) are summarized in Fig. 4. We obtain the following observations based on the results. (i) GFN-based goal-conditioned approaches consistently outperform RL-based goal-conditioned methods (DQN w/ HER), as the latter can easily get trapped in local optima due to its greedy policy. The results validate the promise of training goal-conditioned policies using GFlowNets and pave the way for further advancements in goal-conditioned learning with GFlowNets. (ii) Moreover, our proposed RBS-GFN method significantly outperforms GFN w/ HER and also the stronger OC-GAFN method in terms of sample efficiency and outcome-reaching ability, particularly in larger spaces. In contrast, the performance of GFN w/ HER deteriorates as the complexity of the environment increases, which highlights its limitations in handling large state spaces. (iii) The inferior performance of our method without RBS highlights the significance of our proposed approach. We remark that the superior performance of RBS is attributed to its enhancement of training data with higher quality and diversity.
@@ 5.1.2 Generalization
We now evaluate the generalization ability of our RBS-GFN method to unseen goals and environments, which is important for real-world applications where the agent can encounter novel situations. To evaluate its ability to generalize to unseen goals, we mask $n$ goals from various locations in the map and test the success rates of reaching these unseen goals after the training process, as illustrated in Fig. 6(a) ($n=20$). As shown in Fig. 6(b), RBS-GFN obtains an almost $100$% success rate demonstrating its capacity to effectively determine the required actions to reach novel goals. Moreover, it outperforms our strongest baseline OC-GAFN by an approximately $15\%$ success rate.
@@
Figure 6: (a) DQN fails to generalize to unseen maps with obstacles. (b) RBS-GFN can find diverse trajectories. (c) The average success rate over 200 trials of reaching the goal Refer to caption.
We further investigate its generalization capability to unseen environments. We introduce unseen obstacles during the testing phase following (Kumar et al., 2020), which creates novel environments that the agent has not encountered during training. As shown in Fig. 6(c), RBS-GFN maintains a success rate of almost $100\%$, while OC-GFN obtains a success rate of $9\%$ and the RL-based method DQN completely fails. More unseen maps and corresponding results are provided in Appendix D. The superior performance of RBS-GFN in unseen environments can be attributed to its ability to efficiently discover diverse paths to reach the goal as shown in Fig. 6(b). Although OC-GAFN also has the potential to discover diver paths, its performance is limited by the available training budget. On the other hand, DQN is limited to discovering a single trajectory to reach the goal as shown in Fig. 6(a), making it highly susceptible to failure when the learned path is blocked by unseen obstacles.
@@ 5.1.3 Versatility
In this section, we demonstrate the generality of our approach by integrating it with another recent GFlowNets method based on SubTB (Madan et al., 2023b), whose learning objective is based on $\mathcal{L}_{\rm SubTB}$ as introduced in Eq. (2). We evaluate the goal-reaching performance of RBS-GFN (SubTB) in terms of the success rate in the GridWord task (with $H=10$). As shown in Fig 4, RBS-GFN can also be successfully built upon SubTB with a success rate of $100\%$, and achieves consistent performance gains.
@@ 5.1.4 Ablation Study
In this section, we conduct an in-depth analysis of the key components of RBS-GFN to better understand their effect with a focus on two critical techniques including backward policy regularization and age-based sampling, while we defer the discussion of intensified reward feedback which is essential for scaling up to high-dimensional problems in Appendix A.1.
The backward sampling policy $P_{B}$ plays an important role in synthesizing helpful trajectories for training GC-GFlowNets. We evaluate the effect of different choices of $P_{B}$, including the regularized $P_{B}$ (based on Eq. (6)), a learned backward policy without constraints, and a fixed and uniform one. We compute the entropy of the forward policy $P_{F}$ to measure the ability to generate diverse trajectories of GC-GFlowNets. We further visualize the trajectory distribution in the replay buffer for different choices of $P_{B}$ with t-SNE (Van der Maaten and Hinton, 2008) in Fig. 7(b).

### v2

@@ 3.2 Proposed Method
In this section, we propose a novel approach, Retrospective Backward Synthesis (RBS), which is a simple yet effective method to efficiently tackle these challenges mentioned above in GC-GFlowNets.
@@
Consider a trajectory $\tau=\{s_{0}\to\cdots\to s_{i}\to\cdots\to s_{n}\}$ collected by the forward policy $P_{F}$ of GC-GFlowNets that fails to reach the goal ($s_{n}\neq y$) and receives a zero reward. As illustrated in Fig. 2, RBS utilizes the potential of the backward policy $P_{B}$ to synthesize a backward trajectory $\tau^{\prime}=\{y\to\cdots\to s_{i}^{\prime}\to\cdots\to s_{0}\}$ from the commanded goal. When employing $\tau^{\prime}$ for training, we reverse it to guarantee $\tau^{\prime}$ starts from the initial state $s_{0}$, similar to $\tau$. Therefore, $\tau^{\prime}$ provides a successful training experience as it achieves the goal state, thus enriching training data with positive feedback for GC-GFlowNets to mitigate the sparse reward problem. Unlike previous reward relabeling techniques, such as HER (Andrychowicz et al., 2017), which simply replaces the original goal $y$ with the achieved state $s_{n}$, RBS provides more diverse and informative new training data in the trajectory level. RBS imagines a totally new trajectory $\tau^{\prime}\neq\tau$, thus leading to significant data augmentation and more sample-efficient learning. In practice, we store both collected trajectories $\{\tau_{i}\}_{i=1}^{m}$ and experiences from RBS $\{\tau^{{}^{\prime}}_{i}\}_{i=1}^{m}$ in a replay buffer, and jointly replay the two types of trajectories to optimize GC-GFlowNets.
However, training GC-GFlowNets with RBS may still face the risks of learning instabilities and mode collapse when scaling to longer-horizon and more complex tasks, which may be caused by ineffective reward gradient backpropagation and inefficient experience replay. To address these challenges and further enhance the diversity of generated backward trajectories, we introduce the techniques we have developed in the next paragraphs, where the overall algorithm is summarized in Alg. 1.
@@
3:     Sample a goal $y\sim\mathcal{Y}$ randomly
@@
11:  end for
Age-Based Sampling. Trajectories collected by GC-GFlowNets and the retrospective backward synthesized experiences are stored in a reply buffer for better data utilization. However, uniform sampling of these goal-reaching trajectories often fails to expose the agent to sufficiently diverse experiences that match its evolving learning ability (Fang et al., 2019), which highlights the importance of sampling strategy (Kloek & van Dijk, 1976; Schaul et al., 2016). To guarantee that all experiences are fully considered during training, we introduce an age-based sampling technique. Specifically, age-based sampling assigns the highest priority $p_{\rm max}>0$ to newly added experiences and updates their priority to zero after being learned. Consequently, newly added experiences can be replayed first, while learned experiences are randomly sampled. This prioritization scheme ensures that experiences are leveraged more thoroughly, which balances the exploration of fresh experiences and the exploitation of acquired knowledge.
Backward Policy Regularization. The backward policy can be chosen freely as studied in (Bengio et al., 2023). In the extreme case where $P_{B}$ is set to be a uniform policy, the optimization of GC-GFlowNets becomes challenging as it is not learnable and the data is excessively diverse. On the other hand, specifying $P_{B}$ as a deterministic policy can limit data diversity. We therefore learn the backward policy $P_{B}(\cdot|s^{\prime},\theta)$ within GC-GFlowNets based on the flow consistency criterion in §2.2, instead of specifying it to be a fixed policy, which offers smooth sampling aligned with the current model’s capacity. Yet, when $P_{B}$ degenerates into a deterministic policy, it fails to provide diverse backward trajectories, which can limit the potential of GC-GFlowNets to generalize well. To strike a balance between these two extremes and further enhance the diversity of the imagined trajectories, we introduce a backward policy regularization for $P_{B}$. This regularization term penalizes the Kullback-Leibler (KL) divergence between $P_{B}(\cdot|s^{\prime},\theta)$ and a uniform distribution $\mathcal{U}$, thus encouraging $P_{B}$ to resemble a uniform distribution, while allowing for learning and adaptation. Our training objective can be written as in Eq. (6), where $\gamma$ is the regularization coefficient.
@@
Intensified Reward Feedback. For long-horizon and high-dimensional tasks, a critical factor that affects learning effectiveness is the efficient propagation of the reward signal, which may require a number of steps and affect the learning of intermediate steps. The recent OC-GAFN (Pan et al., 2023a) approach considers the terminal reward at each step (due to the binary nature of rewards), but may introduce stochasticity of the learning signal particularly in the case of the more challenging graph-structured DAG. We propose an efficient technique that intensifies the learning signal, defined as $\mathcal{L}_{\rm GC-GFN}=\left(\log\left[{F_{\theta}(s|y)P_{F}(s^{\prime}|s,y,\theta)}\right]-\log\left[{CR(x,y)P_{B}(s|s^{\prime},y,\theta)}\right]\right)^{2}$ for terminal states $s^{\prime}$, where $C$ is a intensification coefficient to scale the effect of $R(x,y)$. The mechanism behind this technique is that a larger value of $C$ indeed amplifies the gradient of $P_{B}$ by $\log(CR(x,y))$ for terminal states (the detailed derivation can be found in Appendix A). By the intensified reward feedback with a large value of $C$, we effectively strengthen the learning signal propagated backward through the trajectory, enabling more efficient learning and faster convergence in complex environments.
@@
Empirical Validation Fig. 1 compares RBS and OC-GAFN (with HER (Andrychowicz et al., 2017)) in the set generation task with increasing set sizes. The result demonstrates that our RBS approach can scale up to problems with large set sizes, while the previous SOTA method OC-GAFN fails to maintain good performance as the problem complexity increases.
4 Related Work
Generative Flow Networks (GFlowNets). Recently, there have been a number of efforts applying GFlowNets to different important cases, e.g., biological sequence design (Jain et al., 2022), molecule generation (Bengio et al., 2021), combinatorial optimization (Zhang et al., 2023a; 2024), Bayesian structure learning (Deleu et al., 2022). There have also been many works investigating how to improve the training of GFlowNets, enabling them to achieve more efficient credit assignment (Pan et al., 2023b), better exploration (Pan et al., 2023c; Lau et al., 2024), or more effective learning objectives (Bengio et al., 2023; Madan et al., 2023a) that can better handle computational complexity (Bengio et al., 2021), and generalize to stochastic environments (Pan et al., 2023d; Zhang et al., 2023b). GC-GFlowNets learn flows and policies conditioning on outcomes (goals) for reaching any targeted outcomes (Pan et al., 2023a). However, little attention has been given to this topic, leaving this promising direction largely unexplored. Meanwhile, it is challenging to train goal-conditioned policies due to sparse rewards. Our work not only provides a formal definition of GC-GFlowNets but also proposes a novel method called retrospective backward synthesis to significantly improve their training efficiency and success rates.
Goal-Conditioned Reinforcement Learning. Our formulation of goal-conditioned GFlowNets is heavily inspired by the works of goal-conditioned RL. Standard Reinforcement Learning (RL) only requires the agent to finish one specific task defined by the reward function (Schaul et al., 2015), while goal-conditioned RL trains an agent to achieve arbitrary goals as the task specifies (Andrychowicz et al., 2017). Goal-Conditioned RL augments the observation with an additional goal that the agent is required to achieve (Liu et al., 2022). The reward function is usually defined as a binary bonus of reaching the goal. To overcome the challenge of the sparsity of reward function, prior work in goal-conditioned RL has introduced algorithms based on a variety of techniques, such as hindsight relabeling (Andrychowicz et al., 2017; Fang et al., 2018; Yang et al., 2022; Fang et al., 2019; Ding et al., 2019), contrastive learning (Eysenbach et al., 2020; 2022), state-occupancy matching (Durugkar et al., 2021; Ma et al., 2022b) and hierarchical sub-goal planning (Chane-Sane et al., 2021; Kim et al., 2021; Nasiriany et al., 2019). Our work is closely related to hindsight relabeling, denoted as HER (Andrychowicz et al., 2017), which relabels any experience with some commanded goal to the goal that was actually achieved in order to learn from failures. HER can generate non-negative rewards to alleviate the negative sparse reward problem, even if the agent did not complete the task. However, the agent using HER still suffers from low sample efficiency on large-scale problems due to its limitation in operating only on the observed trajectories, while our method can imagine new trajectories with positive rewards for policy training.
@@ 5 Experiments
In this section, we conduct extensive experiments to investigate our Retrospective Backward Synthesis (RBS) method to answer the following key questions: i) How does RBS-GFN compare against previous baselines in terms of sample efficiency and success rates? ii) Can RBS-GFN scale to complex and high-dimensional environments? iii) Can RBS-GFN effectively generalize to unseen goals and unseen environments? iv) Is RBS-GFN general and can be built upon different GFlowNets training objectives? v) What are the effects of important components in RBS-GFN?
@@ 5.1 GridWorld
We first conduct a series of experiments based on the GridWorld environment (Bengio et al., 2021), in which the model learns to achieve any given goals starting in a $H\times H$ grid. Specifically, we investigate mazes with increasing horizons $H$ ($32$, $64$, and $128$), respectively, resulting in different levels of difficulty categorized as small, medium, and large.
We compare our proposed RBS-GFN approach with the following state-of-the-art baselines. (i) GFN w/ HER (Andrychowicz et al., 2017) is a GC-GFlowNets that relabels the negative reward in a failed trajectory with a positive reward. (ii) OC-GAFN (Pan et al., 2023a) is a recent method that utilizes contrastive learning to complement successful experiences, and employs a trained Generative Augmented Flow Network (GAFN; Pan et al. (2023c)) as an exploratory component to generate diverse outcomes $y$, which are subsequently provided to sample goal-conditioned trajectories. (iii) DQN w/ HER leverages both deep Q-learning algorithm (Mnih et al., 2013; Jang et al., 2019) and HER technique (Andrychowicz et al., 2017) to learn a near-optimal policy. This baseline is used to ablate the effects of GFlowNet-based training compared with RL-style methods. To ensure fairness, each baseline has the same model architecture and training steps as RBS-GFN, and we follow the experimental setup for hyperparameters as in (Pan et al., 2023a). We run each algorithm with three different seeds and report their performance in mean and standard deviation. A more detailed description of the experimental setup can be found in Appendix C.
@@ 5.1.1 Performance Comparison
The success rates for different methods for increasing sizes of the GridWorld environment (including small, medium, and large) are summarized in Fig. 4. We obtain the following observations based on the results. (i) GFN-based goal-conditioned approaches consistently outperform RL-based goal-conditioned methods (DQN w/ HER), as the latter can easily get trapped in local optima due to its greedy policy. The results validate the promise of training goal-conditioned policies using GFlowNets and pave the way for further advancements in goal-conditioned learning with GFlowNets. (ii) Moreover, our proposed RBS-GFN method significantly outperforms GFN w/ HER and is stronger than the OC-GAFN method in terms of sample efficiency and outcome-reaching ability, particularly in larger spaces. In contrast, the performance of GFN w/ HER deteriorates as the complexity of the environment increases, which highlights its limitations in handling large state spaces. (iii) The inferior performance of our method without RBS highlights the significance of our proposed approach. We remark that the superior performance of RBS is attributed to its enhancement of training data with higher quality and diversity.
@@
5.1.2 Special Case: Offline GC-GFlowNets
Given the potential of learning general goal-reaching policies solely from given offline datasets Ma et al. (2022a), we investigate the offline goal-conditioned scenario where GC-GFlowNets learn from fixed offline data without interacting with the environments. As a result, all the baselines are restricted in limited training datasets, while RBS-GFN can synthesize a number of new trajectories to enhance the learning process. We compare our method with the two strongest baselines, OC-GAFN and vanilla goal-conditioned GFN, and evaluate them on three different sizes of GridWorld. As demonstrated in Fig. 5, our method achieves nearly 100% success rates across all scenarios, while the performance of the baselines declines significantly as the problem size increases. The experimental details and learning curves can be found in Appendix C.3.
@@ 5.1.3 Generalization
Figure 5: Success rates on GridWorld tasks with different sizes.
We now evaluate the generalization ability of our RBS-GFN method to unseen goals and environments, which is important for real-world applications where the agent can encounter novel situations. To evaluate its ability to generalize to unseen goals, we mask $n$ goals from various locations in the map and test the success rates of reaching these unseen goals after the training process, as illustrated in Fig. 7(a) ($n=20$). As shown in Fig. 7(b), RBS-GFN obtains an almost $100$% success rate, demonstrating its capacity to effectively determine the required actions to reach novel goals. Moreover, it outperforms our strongest baseline OC-GAFN by an approximately $15\%$ success rate.
@@
Figure 7: (a) DQN fails to generalize to unseen maps with obstacles. (b) RBS-GFN can find diverse trajectories. (c) The average success rate over 200 trials of reaching the goal Refer to caption
We further investigate its generalization capability to unseen environments. We introduce unseen obstacles during the testing phase following (Kumar et al., 2020), which creates novel environments that the agent has not encountered during training. As shown in Fig. 7(c), RBS-GFN maintains a success rate of almost $100\%$, while OC-GFN obtains a success rate of $9\%$ and the RL-based method DQN completely fails. More unseen maps and corresponding results are provided in Appendix D. The superior performance of RBS-GFN in unseen environments can be attributed to its ability to efficiently discover diverse paths to reach the goal as shown in Fig. 7(b). Although OC-GAFN also has the potential to discover diver paths, its performance is limited by the available training budget (i.e., fewer pre-training iterations). On the other hand, DQN is limited to discovering a single trajectory to reach the goal as shown in Fig. 7(a), making it highly susceptible to failure when the learned path is blocked by unseen obstacles.
@@ 5.1.4 Versatility
In this section, we demonstrate the generality of our approach by integrating it with another recent GFlowNets method based on SubTB (Madan et al., 2023b), whose learning objective is based on $\mathcal{L}_{\rm SubTB}$ as introduced in Eq. (2). We evaluate the goal-reaching performance of RBS-GFN (SubTB) in terms of the success rate in the GridWord task (with $H=10$).
As shown in Fig 4, RBS-GFN can also be successfully built upon SubTB with a success rate of $100\%$, and achieves consistent performance gains.
@@ 5.1.5 Ablation Study
In this section, we conduct an in-depth analysis of the key components of RBS-GFN to better understand their effect with a focus on two critical techniques, including backward policy regularization and age-based sampling, while we defer the discussion of intensified reward feedback, which is essential for scaling up to high-dimensional problems in Appendix A.1.
The backward sampling policy $P_{B}$ plays an important role in synthesizing helpful trajectories for training GC-GFlowNets. We evaluate the effect of different choices of $P_{B}$, including the regularized $P_{B}$ (based on Eq. (6)), a learned backward policy without constraints, and a fixed and uniform one. We compute the entropy of the forward policy $P_{F}$ to measure the ability to generate diverse trajectories of GC-GFlowNets. We further visualize the trajectory distribution in the replay buffer for different choices of $P_{B}$ with t-SNE (Van der Maaten & Hinton, 2008) in Fig. 8(b).